# Breaking the Bulkhead: Demystifying Cross-Namespace Reference Vulnerabilities in Kubernetes Operators

### v1

@@ 4.2. Insecure Namespace-Scoped Resource Reference
The Operator then processes the newly created malicious resource. The Operator extracts the fields in the malicious resource, operates the victim resource located in the victim namespace, and inadvertently leaks or tampers with sensitive information. Thus, the attacker effectively escalates their privileges, bypassing Kubernetes’ namespace isolation, gaining unauthorized access to resources that should have remained secure.
Example. A common real-world scenario occurs when an Operator manages applications consuming credentials (e.g., API Secret Key) stored in Kubernetes Secrets. As per Kubernetes official security practice (Goo, 2024), Secrets should only be referenced strictly within the same namespace to maintain proper isolation.
Figure 3. Insecure Namespace-Scoped Resource Reference Sample
[…]
Insecure Namespace-Scoped Resource Reference Sample
@@
Impact. Insecure Namespace-Scoped Resource Reference vulnerabilities fundamentally enable attackers to escalate privileges by allowing them to reference and manipulate resources in namespaces beyond their legitimate access. Further impact of this vulnerability heavily depends on how the Operator processes and utilizes the referenced resources, as well as the nature of the referenced resources themselves. For instance, if the referenced resource is a Kubernetes Secret containing sensitive credentials like API Tokens, an attacker may obtain unauthorized access to applications, databases, or cloud infrastructure. If the Operator not only reads but also modifies referenced resources, attackers might disrupt service availability, modify application configurations, or inject malicious workloads. Thus, the severity and scope of the impact are highly context-dependent, ranging from sensitive information leakage to complete cluster compromise, based on the type and usage of the improperly referenced resource.
4.3. Insecure Cluster-Scoped Resource Reference
Insecure Cluster-Scoped Resource Reference occurs when a Kubernetes Operator processes a namespace-scoped resource and interacts with cluster-scoped resources. Unlike namespace-scoped resources that remain isolated within specific namespaces, cluster-scoped resources affect the entire Kubernetes cluster. If an Operator allows users to influence these cluster-scoped resources through namespace-scoped resources, it creates a pathway for attackers to escalate privileges and potentially compromise the entire cluster.
@@ 4.3. Insecure Cluster-Scoped Resource Reference
The basic attack workflow for this vulnerability starts with an attacker creating a malicious resource in their authorized namespace, whose fields are leveraged by the Operator to reference a cluster-scoped resource. The Operator, running with elevated cluster-level privileges, processes this malicious input and subsequently performs operations on the referenced cluster-scoped resource. As cluster-scoped resources inherently affect the entire Kubernetes environment, these unauthorized accesses and manipulations enable attackers to escalate privileges beyond their initial namespace boundaries and gain control or influence over all the other namespaces.
@@
Figure 4. Insecure Cluster-Scoped Resource Reference Sample
[…]
Insecure Cluster-Scoped Resource Reference Sample
Operators often interact with cluster-scoped resources, ClusterRole or ClusterRoleBinding, because some applications require elevated or cluster-wide permissions to operate correctly. An insecure implementation occurs when an Operator accepts namespace-scoped resources and assigns a ClusterRole and ClusterRoleBinding to the requesting namespace. An attacker restricted in a namespace could thus craft the malicious resource illustrated in the YAML file of Figure 4. He deploys a resource with metadata.namespace setting to attacker, which means the resource is deployed in the attacker namespace.
The Operator monitors App custom resources. It finds the malicious resource deployed by the attacker, and creates ClusterRoleBinding towards a ServiceAccount in the attacker’s namespace. Thus, the attacker can impersonate the ServiceAccount in his authorized namespace to escalate his privilege, gaining cluster-level permissions.
Impact. Insecure Cluster-Scoped Resource References allow attackers to escalate privileges and affect resources across all namespaces in a cluster. The specific severity and effect of this vulnerability depend on which cluster-scoped resources the Operator interacts with. For instance, if an Operator insecurely assigns a ClusterRole or ClusterRoleBinding as dictated by namespace-scoped resources, an attacker can gain cluster-level permissions. The detail permissions assigned depend on the implementation of Operators.
5. Cross-Namespace In The Wild
To assess the prevalence of the vulnerabilities described in Section 4, we conducted a large-scale measurement of real-world Kubernetes Operators and disclosed our findings to affected vendors.
5.1. Measurement Methodology
5.1.1. Overview
To understand how widespread the vulnerabilities are in real-world Kubernetes Operators, we perform a systematic measurement illustrated in Figure 5, which consisting of the following steps:
Figure 5. Measurement Workflow
  1. (1)
@@ 5.1.1. Overview
  2. (2)
    Resource Type Identification: Resource types (either Kubernetes built-in resources or custom resources) used by each Operator are extracted, and their declared scopes (either namespace-scoped or cluster-scoped) are identified based on the code.
  3. (3)
    Vulnerability Detection: The analysis identifies whether Operators process namespace-scoped resources but conduct insecure cross-namespace reference behavior, as depicted in Section 4.2 and Section 4.3.
  4. (4)
@@
Since the two dominant Kubernetes Operator frameworks with the highest GitHub Stars, Kubebuilder (kub, 2025c) and Operator SDK (ope, 2025), are written in Golang, the collection specifically targets Operators implemented in Golang. We adopted CodeQL (Cod, 2025) v2.17.4 to analyze Operators. The entire CodeQL query suite uses around 1,500 lines of QL rules.
Table 1. Common Operator-Related Libraries
@@
k8s.io/api (k8s, 2025) K8s Built-In Resource Specifications
k8s.io/apimachinery (api, 2025) K8s Metadata Specifications
client-go (cli, 2025) K8s Official Client
client-gen (gen, 2025) K8s Official Client Generator
controller-runtime (con, 2025) Controller Client
Kubebuilder (kub, 2025c) Operator Framework
Operator SDK (ope, 2025) Operator Framework
@@
To enhance the measurement process, 7 commonly used libraries listed in Table 1 were modeled to accurately resolve and track Kubernetes interactions within collected Operator implementations. They contain specifications of native Kubernetes resources, namespace-related data structures, and functions for Operators to manipulate resource. The detail is elaborated later.
5.1.2. Operator Collection
@@
The dataset of Kubernetes Operators analyzed was collected by crawling GitHub repositories. To achieve this, GitHub Search API (git, 2022) was utilized with the query string ”Kubernetes Operator language:go”. The collection process strictly adhered to GitHub’s API rate limits and usage guidelines to responsibly retrieve relevant Operator repositories.
@@
After collecting Operators from GitHub, we set up CodeQL databases for each Operator. Operators that cannot be compiled to generate the CodeQL database due to errors, like syntax and dependency errors, are eliminated, and the final set contains 2,268 Operators.
5.1.3. Resource Type Identification
@@
The first critical step in detecting vulnerabilities is Resource Type Identification, as the attack requires the attacker to initiate operations using a namespace-scoped resource they are authorized to create in their own namespace. The analysis separately handles Kubernetes Built-in Resources and Operator-defined Custom Resources.
@@
Custom Resource Identification. For Custom Resources defined by the Operators, their data structures can be explicitly extracted from the source code. In Kubernetes, each resource structure must contain a field of type TypeMeta (defined by the Apimachinery library (api, 2025)), which acts as the unique identifier of a resource type. Thus the analyzer extracts all struct types in Operators and filter those with TypeMeta fields. This outputs all custom resource types in Operators.
@@
To further identify the scope of each resource type (namespace or cluster), common Operator frameworks like Kubebuilder (kub, 2025c) and Operator SDK (ope, 2025), as well as Kubernetes’ official client code generator (gen, 2025), require developers to explicitly decorate cluster-scoped resource structs using special marker annotations ”+genclient:nonNamespaced” or ”+kubebuilder:resource:scope:Cluster”. By detecting these markers in the Custom Resource struct definitions, the analyzer reliably identifies Custom Resource types in Operators as well as their scopes.
@@
Kubernetes Built-in Resources. Unlike Custom Resources, the built-in Kubernetes resource specifications are imported from the external k8s.io/api (k8s, 2025) library to Operators, thus their source code and scope markers are not directly accessible for CodeQL. Therefore, the analyzer adopts an alternative method. Specifically, Operators would ultimately depend on the client-go library (cli, 2025). Each type of built-in Kubernetes resource is uniquely associated with a typed client provided by the client-go library (cli, 2025). Each typed client is constructed by methods in its corresponding Getter interfaces under the k8s.io/client-go/kubernetes/typed package. For example, considering the built-in resource type Pod, there is a uniquely associated Pod client. The Pod client is constructed by the only method in the PodGetter interface. By extracting the return types of methods in all Getter interfaces, the analyzer identifies all built-in types and their clients.
@@
To determine the scope of built-in resources, the analysis leverages the only method in the Getter interface of each typed client. Specifically, namespace-scoped resource clients in client-go require a namespace parameter in their constructor to specify the target namespace. In contrast, constructors for cluster-scoped resource clients do not require such a namespace argument. By counting and verifying constructor parameters, the analysis distinguishes namespace-scoped from cluster-scoped built-in resources.
5.1.4. Vulnerability Detection
@@
This step determines whether an attacker-controlled input can influence sensitive operations that cross namespace boundaries or impact the whole cluster. To achieve this, the analysis uses interprocedural taint tracking, tracing the propagation of data originating from the namespace-scoped resource objects to insecure reference sites in the controller logic.
@@
Table 2. Resource Namespace Setters
@@
Thus, the analysis tracks data flow from namespace-scoped resource objects to namespace setters (listed in Table 2) that are leveraged to specify the namespace of a resource. We identify 3 struct fields that can store namespace values in 2 libraries. We identify 4 types of functions in 3 libraries that can be used to set the namespace field of a resource object or set up a typed client towards a specific namespace.
It is worth noting that the 3 fields (listed in Table 2) of namespace-scoped resource objects are excluded from the taint source, as these fields denote the namespace where this resource is deployed. Since the attackers are only authorized to access their namespace, the namespace fields of attacker-controlled resources are always the attacker-authorized namespace. If these fields sink in the referenced resources’ namespace fields, it means the referenced resources are also in the attacker-authorized namespace. Thus, no cross-namespace operation is conducted.
@@
Insecure Cluster-Scoped Resource Reference. This analysis aims to detect whether attacker-controlled input can influence cluster-scoped resource. Since these cluster-scoped resources affect the entire Kubernetes cluster, any modification to them based on namespace-scoped input represents a significant privilege escalation risk. Thus, the analysis tracks data flow from the identified namespace-scoped resource objects into any cluster-scoped resource objects. If tainted input is used to construct such cluster-scoped resource objects, the Operator is flagged as vulnerable to insecure cluster-scoped resource references.
5.1.5. Summary
To understand what an adversary can do to which kind of resource, the measurement further identifies insecure referenced resource types and operations towards these resources.
@@ 5.1.5. Summary
For insecure namespace-scoped resource references, the affected resource type identification depends on the type of sink site in the previous step. If the previous data flow sinks at the three fields, WithNamespace(), or SetNamespace() methods, then the analyzer further tracks interprocedural data flow from the previous sinks to any namespace-scoped resource objects to identify affected resource objects and types. If the previous data flow sinks at the constructor of a typed client, then the resource type is the one associated with that typed client.
@@
5.2. Measurement Result
We conducted measurements on 2,268 Operators crawled from GitHub to assess the real-world impacts of insecure cross-namespace references and answer the following research questions:
@@ 5.2. Measurement Result
  • RQ3: What can an attacker do towards cross-namespace referenced resources?
@@
5.2.1. RQ1: How Many Operators Are Potentially Vulnerable To Cross-Namespace Reference?
Figure 6. Percentage of Affected Operators
To assess the prevalence of insecure cross-namespace reference vulnerabilities, we analyzed a dataset of 2,268 real-world Kubernetes Operators collected from GitHub. Each Operator was examined to determine whether attacker-controlled namespace-scoped resources can influence operations across namespace boundaries. The results are summarized as follows:
  • 143 Operators (6.3%) only include insecure namespace-scoped resources references that can specify or influence operations on other namespaces.
  • 122 Operators (5.4%) only contain insecure cluster-scoped resource references, where attacker-controlled namespace-scoped resources can affect cluster-scoped resources.
  • 52 Operators (2.3%) allow both types of references, posing risks of privilege escalation at both the namespace and cluster level.
These findings show that a non-negligible portion (over 14%) includes logic that may lead to privilege escalation, highlighting a widespread but largely overlooked security concern in the Kubernetes ecosystem.
5.2.2. RQ2: What Resources Can Be Cross-Namespace Referenced By Attackers?
Table 3. Major Insecurely Referenced Resource Type

### v2

@@ IV-B Insecure Namespace-Scoped Resource Reference
The Operator then processes the created malicious resource. The Operator extracts the fields in the malicious resource, operates the victim resource located in the victim namespace, and inadvertently leaks or tampers with sensitive information. Thus, the attacker effectively escalates their privileges, bypassing Kubernetes’ namespace isolation, gaining unauthorized access to resources that should have remained secure.
Example. A common real-world scenario occurs when an Operator manages applications consuming credentials (e.g., API Secret Key) stored in Kubernetes Secrets. As per Kubernetes official security practice [Goodprac92:online], Secrets should only be referenced strictly within the same namespace to maintain proper isolation.
Figure 3: Insecure Namespace-Scoped Resource Reference Sample
@@
Impact. Insecure NS-Scoped Resource Reference vulnerabilities fundamentally enable attackers to escalate privileges by allowing them to reference and manipulate resources in namespaces beyond their legitimate access. Further impact of this vulnerability heavily depends on how the Operator processes and utilizes the referenced resources, as well as the nature of the referenced resources themselves. For instance, if the referenced resource is a Kubernetes Secret containing sensitive credentials like API Tokens, an attacker may obtain unauthorized access to applications, databases, or cloud infrastructure. If the Operator not only reads but also modifies referenced resources, attackers might disrupt service availability, modify application configurations, or inject malicious workloads. Thus, the severity and scope of the impact are highly context-dependent, ranging from sensitive information leakage to complete cluster compromise, based on the type and usage of the improperly referenced resource.
IV-C Insecure Cluster-Scoped Resource Reference
Insecure Cluster-Scoped Resource Reference occurs when a Kubernetes Operator processes a ns-scoped resource and interacts with cluster-scoped resources. Unlike ns-scoped resources that remain isolated within specific namespaces, cluster-scoped resources affect the entire Kubernetes cluster. If an Operator allows users to influence these cluster-scoped resources through ns-scoped resources, it creates a pathway for attackers to escalate privileges and potentially compromise the entire cluster.
@@ IV-C Insecure Cluster-Scoped Resource Reference
The basic attack workflow for this vulnerability starts with an attacker creating a malicious resource in their authorized namespace, whose fields are leveraged by the Operator to reference a cluster-scoped resource. Kubernetes RBAC accepts the deployment of the malicious resource because it only knows that the attacker has created a resource under their authorized namespace, but does not know if the resource leads to privilege escalation. The Operator, running with elevated cluster-level privileges, processes this malicious input and subsequently performs operations on the referenced cluster-scoped resource. As cluster-scoped resources inherently affect the entire Kubernetes environment, these unauthorized accesses and manipulations enable attackers to escalate privileges beyond their initial namespace boundaries and gain control or influence over all the other namespaces.
@@
Figure 4: Insecure Cluster-Scoped Resource Reference Sample
ClusterRole and ClusterRoleBinding are two cluster-scoped built-in resources in Kubernetes. It can grant cluster-wide permissions to a ServiceAccount, a built-in namespace-scoped resource that provides an identity for applications to access the Kubernetes API. Some Kubernetes applications require mounting a ServiceAccount with cluster-wide permissions to access the Kubernetes API and operate correctly. Thus, when deploying such an application in a specified namespace, their Operators will create a ServiceAccount in that namespace, mount it on the application, and create a ClusterRoleBinding to grant cluster-wide permission to that ServiceAccount. An insecure implementation occurs when an Operator accepts ns-scoped resources and creates a ClusterRoleBinding to a ServiceAccount in the requesting namespace. An attacker restricted to a namespace could thus craft the malicious resource illustrated in the YAML file of Figure 4. He deploys a resource with metadata.namespace setting to attacker, which means the resource is deployed in the attacker namespace.
The Operator monitors App custom resources. It finds the malicious resource deployed by the attacker, creates a ServiceAccount in the attacker’s namespace, and then creates a ClusterRoleBinding towards that ServiceAccount to grant it cluster-wide permissions. Since the ServiceAccount is created in the attacker’s namespace, the attacker can impersonate the ServiceAccount to escalate his privilege, gaining cluster-wide permissions granted by ClusterRoleBinding. In short, such a vulnerability can be leveraged to directly elevate the permissions of attackers.
Impact. Insecure Cluster-Scoped Resource References allow attackers to escalate privileges and affect resources across all namespaces in a cluster. The specific severity and effect of this vulnerability depend on which cluster-scoped resources the Operator interacts with. For instance, if an Operator insecurely creates a ClusterRole or ClusterRoleBinding as dictated by ns-scoped resources, an attacker can gain cluster-level permissions. The detailed permissions assigned depend on the implementation of Operators.
V Cross-Namespace In The Wild
To assess the prevalence of the vulnerabilities described in Section IV, we conducted a large-scale measurement of real-world Kubernetes Operators and disclosed our findings to affected vendors.
V-A Measurement Methodology
V-A1 Overview
To understand how widespread the vulnerabilities are in real-world Kubernetes Operators, we perform a systematic measurement illustrated in Figure 5, which consists of the following steps:
Figure 5: Measurement Workflow
  1. 1.
@@ V-A1 Overview
  2. 2.
    Resource Type Identification: Resource types (either Kubernetes built-in resources or custom resources) used by each Operator are extracted, and their declared scopes (either ns-scoped or cluster-scoped) are identified based on the code.
  3. 3.
    Vulnerability Detection: The analysis identifies whether Operators process ns-scoped resources but conduct insecure cross-ns reference behavior, as depicted in Section IV-B and Section IV-C.
  4. 4.
@@
Among all Kubernetes development frameworks recommended by the Kubernetes official [operator], Golang frameworks (i.e. Kubebuilder [kubebuilder] and Operator SDK [operatorsdk]) own the highest GitHub Stars and dominates with around 2.4k Operators, followed by Python (331 Operators), Shell (167 Operators), Java (134 Operators), Rust (96 Operators), and .NET (29 Operators). Given the overwhelming amount of Golang-based Operators, the collection specifically targets Operators implemented in Golang. We adopted CodeQL [CodeQL] v2.17.4 to analyze Operators. The entire CodeQL query suite uses around 1,500 lines of QL rules, covering detailed modeling of major Kubernetes Go libraries, enabling analysis of Kubernetes applications beyond Operators.
TABLE I: Common Operator-Related Libraries
@@
k8s.io/api [k8sapi] K8s Built-In Resource Specifications
k8s.io/apimachinery [apimachinery] K8s Metadata Specifications
client-go [clientgo] K8s Official Client
client-gen [gen] K8s Official Client Generator
controller-runtime [controllerruntime] Controller Client
@@
To enhance the measurement process, 5 commonly used libraries listed in Table I were modeled to accurately resolve and track Kubernetes interactions within collected Operator implementations. They contain specifications of native Kubernetes resources, namespace-related data structures, and functions for Operators to manipulate resources. The detail is elaborated later.
V-A2 Operator Collection
@@
The dataset of Kubernetes Operators analyzed was collected by crawling GitHub repositories. To achieve this, GitHub Search API [githubsearch] was utilized with the query string “Kubernetes Operator language: go”. The collection process strictly adhered to GitHub’s API usage policies to responsibly retrieve relevant Operator repositories.
@@
After collecting Operators from GitHub, we set up CodeQL databases for each Operator. 13 Operators that cannot be compiled to generate the CodeQL database due to errors, like syntax and dependency errors, are eliminated, and the final set contains 2,268 Operators.
V-A3 Resource Type Identification
@@
The first critical step in detecting vulnerabilities is Resource Type Identification, as the attack requires the attacker to initiate operations using a ns-scoped resource they are authorized to create in their namespace. The analysis separately handles Kubernetes Built-in Resources and Operator-defined Custom Resources.
@@
Custom Resource Identification. For Custom Resources defined by the Operators, their data structures can be explicitly extracted from the source code. In Kubernetes, each resource structure must contain a field of type TypeMeta (defined by the Apimachinery library [apimachinery]), which acts as the unique identifier of a resource type. Thus, the analyzer extracts all struct types in Operators and filters those with TypeMeta fields. This outputs all custom resource types in Operators.
@@
To further identify the scope of each resource type (ns or cluster), common Kubernetes frameworks like Kubebuilder [kubebuilder] and Operator SDK [operatorsdk], as well as Kubernetes’ official client code generator [gen], require developers to explicitly decorate cluster-scoped resource structs using special marker annotations “+genclient:nonNamespaced” or “+kubebuilder:resource:scope:Cluster”. By detecting these markers in the Custom Resource struct definitions, the analyzer reliably identifies Custom Resource types in operators and their scopes.
@@
Kubernetes Built-in Resources. Unlike Custom Resources, the built-in Kubernetes resource specifications are imported from the external k8s.io/api [k8sapi] library to Operators, thus their source code and scope markers are not directly accessible for CodeQL. Therefore, the analyzer adopts an alternative method. Specifically, Operators would ultimately depend on the client-go library [clientgo]. Each type of built-in Kubernetes resource is uniquely associated with a typed client provided by the client-go library [clientgo]. Each typed client is constructed by methods in its corresponding Getter interfaces under the k8s.io/client-go/kubernetes/typed package. For example, considering the built-in resource type Pod, there is a uniquely associated Pod client. The Pod client is constructed by the only method in the PodGetter interface. By extracting the return types of methods in all Getter interfaces, the analyzer identifies all built-in types and their clients.
@@
To determine the scope of built-in resources, the analysis leverages the only method in the Getter interface of each typed client. Specifically, ns-scoped resource clients in client-go require a namespace parameter in their constructor to specify the target namespace. In contrast, constructors for cluster-scoped resource clients do not require such a namespace argument. By counting and verifying constructor parameters, the analysis distinguishes ns-scoped from cluster-scoped built-in resources.
V-A4 Vulnerability Detection
@@
This step determines whether an attacker-controlled input can influence sensitive operations that cross namespace boundaries or impact the whole cluster. To achieve this, the analysis uses interprocedural taint tracking, tracing the propagation of data originating from the ns-scoped resource objects to insecure reference sites in the controller logic.
@@
TABLE II: Resource Namespace Setters
@@
Thus, the analysis tracks data flow from ns-scoped resource objects to namespace setters (listed in Table II) that are leveraged to specify the namespace of a resource. By systematic code reviews of common Operator libraries, we identify 3 struct fields that can store namespace values in 2 libraries and identify 4 types of functions in 3 libraries that can be used to set the namespace field of a resource object or set up a typed client towards a specific namespace.
It is worth noting that the 3 fields (listed in Table II) of ns-scoped resource objects are excluded from the taint source, as these fields denote the namespace where this resource is deployed. Since the attackers are only authorized to access their namespace, the namespace fields of attacker-controlled resources are always the attacker-authorized namespace. If these fields sink in the referenced resources’ namespace fields, it means the referenced resources are also in the attacker-authorized namespace. Thus, no cross-ns operation is conducted.
@@
Insecure Cluster-Scoped Resource Reference. This analysis aims to detect whether attacker-controlled input can influence cluster-scoped resources. Since cluster-scoped resources affect the entire Kubernetes cluster, any modification to them based on ns-scoped input represents a significant privilege escalation risk. Thus, the analysis tracks data flow from the identified ns-scoped resource objects into any cluster-scoped resource objects. If tainted input is used to construct cluster-scoped resource objects, the Operator is flagged as vulnerable to insecure cluster-scoped resource references.
V-A5 Summary
To understand what an adversary can do to which kind of resource, the measurement further identifies insecurely referenced resource types and operations towards these resources.
@@ V-A5 Summary
For insecure ns-scoped resource references, the affected resource type identification depends on the type of sink site in the previous step. If the previous data flow sinks at the three fields, WithNamespace(), or SetNamespace() methods, then the analyzer further tracks interprocedural data flow from the previous sinks to any ns-scoped resource objects to identify affected resource objects and types. If the previous data flow sinks at the constructor of a typed client, then the resource type is the one associated with that typed client.
@@
V-B Measurement Result
We conducted measurements on 2,268 Operators crawled from GitHub to assess the real-world impacts of insecure cross-ns references. The measurement suite was run on a Windows 11 machine with an Intel i7-10700K CPU (3.80GHz) and 32GB RAM. The suite cost 40026 seconds, with 17.6 seconds per Operator on average. In this part, we answered the following research questions:
@@ V-B Measurement Result
  • RQ3: What can attackers do towards cross-namespace referenced resources?
@@
V-B1 RQ1: How Many Operators Are Potentially Vulnerable To Cross-Namespace Reference?
Figure 6: Percentage of Affected Operators
To assess the prevalence of insecure cross-ns reference vulnerabilities, we analyzed a dataset of 2,268 real-world Kubernetes Operators collected from GitHub. Each Operator was examined to determine whether attacker-controlled ns-scoped resources can influence operations across namespace boundaries. The results illustrated that 318 Operators were potentially vulnerable, affecting 282 Operator providers. Specifically:
  • 196 Operators (8.6%) include insecure ns-scoped resource references that can specify or influence operations on other namespaces.
  • 175 Operators (7.7%) contain insecure cluster-scoped resource references, where attacker-controlled ns-scoped resources can affect cluster-scoped resources.
  • 53 Operators (2.3%) allow both types of references, posing risks of privilege escalation at both the namespace and cluster level.
These findings illustrate that a non-negligible portion (over 14%) includes logic that may lead to privilege escalation, highlighting a widespread but largely overlooked security concern in the Kubernetes ecosystem.
TABLE III: Operator Type Summary
Type Validated? # of Flagged # of Validated
Cluster Management 56 11
Database 52 9
Networking 34 6
Cloud Management 25 3
CI/CD 23 3
Metrics 22 3
Storage 17 4
Gadgets 12 3
Secret Management 10 1
Device 9 2
Workflow 8 1
Testing 7 4
Identity × 7 0
AI 7 2
OS Management 6 1
Searching 4 1
Security × 4 0
Framework × 3 0
Config Management 3 1
CMS × 2 0
Financial × 2 0
Registry × 2 0
Backup × 1 0
Collaboration × 1 0
Streaming × 1 0
Total 16 318 55
Table III illustrates the type distribution of flagged Operators. These Operators cover a wide range of usage types in the real world, where Cluster Management, Database, and Networking are most prevalent.
We further assessed the accuracy of our measurement. Due to the lack of existing datasets or detection tools for this new class of vulnerabilities, it is infeasible to validate all Operators at scale manually. Thus, we primarily focused on false positives, evaluating if flagged Operators were truly vulnerable. We randomly reviewed 55 flagged Operators without any specific tailored criteria, covering major types of Operators as listed in Table III. Although this sampling may not cover all potential Operator variations or directly conclude the perfect representative, it provides a practical basis to estimate the accuracy of our detection by covering the major types (or most prevalent types) of Operators, as we shown in Table III. Among the 55 cases, only 5 were false positives. This result indicates that the vast majority of cases flagged are indeed vulnerable, which supports the reliability of our measurement concerning the prevalence of vulnerability. While false negatives are difficult to quantify due to the lack of ground truth, our analysis focused on a range of commonly observed cross-namespace patterns, which may not capture all potential variations but still reflect realistic threats and led to confirmed vulnerabilities. Overall, the measurement provides strong evidence that Cross-NS Reference Vulnerabilities are non-negligible in the real-world ecosystem.
V-B2 RQ2: What Resources Can Be Cross-Namespace Referenced By Attackers?
TABLE IV: Major Insecurely Referenced Resource Type